### v1

# Beyond GPT-5: Making LLMs Cheaper and Better via Performance–Efficiency Optimized Routing

Abstract
Balancing performance and efficiency is a central challenge in large language model (LLM) advancement. GPT-5 addresses this with test-time routing, dynamically assigning queries to either an efficient or a high-capacity model during inference. In this work, we present Avengers-Pro, a test-time routing framework that ensembles LLMs of varying capacities and efficiencies, providing a unified solution for all performance-efficiency tradeoffs. The Avengers-Pro embeds and clusters incoming queries, then routes each to the most suitable model based on a performance-efficiency score. Across 6 challenging benchmarks and 8 leading models—including GPT-5-medium, Gemini-2.5-pro, and Claude-opus-4.1—Avengers-Pro achieves state-of-the-art results: by varying a performance-efficiency trade-off parameter, it can surpass the strongest single model (GPT-5-medium) by +7% in average accuracy. Moreover, it can match the average accuracy of the strongest single model at 27% lower cost, and reach $\sim$90% of that performance at 63% lower cost. Last but not least, it achieves a Pareto frontier, consistently yielding the highest accuracy for any given cost, and the lowest cost for any given accuracy, among all single models. Code is available at https://github.com/ZhangYiqun018/AvengersPro.
Figure 1: Avengers-Pro optimizes the trade-off between performance (accuracy) and efficiency (cost). (A) By varying a trade-off parameter $\alpha$, Avengers-Pro establishes a Pareto frontier. Compared to all single models, it achieves the highest accuracy for any given cost, and achieves the lowest cost for any given accuracy. (B) With comparable cost, Avengers-Pro outperforms the strongest single model GPT-5-medium by $7.1\%$. With comparable performance, Avengers-Pro achieves a $26.9\%$ cost reduction compared to GPT-5-medium.
1 Introduction
A fundamental dilemma in LLM advancement is the trade-off between performance and efficiency. To navigate this, a defining feature of GPT-5 is its test-time routing between models. As described in Introducing GPT-5111https://openai.com/index/introducing-gpt-5/:
“GPT-5 is a unified system with a smart, efficient model that answers most questions, a deeper reasoning model (GPT-5 thinking) for harder problems, and a real-time router that quickly decides which to use based on conversation type, complexity … ”
The efficient model offers lower computational cost and latency at the expense of capability, while the deeper reasoning model incurs higher cost and latency but delivers greater capability. During inference, GPT-5’s router dynamically assigns each query to exactly one model, striking a balance between performance and efficiency.
In this work, we advance test-time routing to optimize the performance–efficiency trade-off. We build upon our earlier work Avengers zhang2025avengers —which showed that a simple routing recipe using ten models ($\sim$7B parameters each) surpass GPT-4.1 and 4.5 across 15 datasets—and introduce the Avengers-Pro. With a focus on performance-efficiency trade-off, the Avengers-Pro operates through three lightweight operations: (i) embedding: encode queries using a text embedding model, (ii) clustering: group queries by semantic similarity, and (iii) scoring: evaluate models within each cluster based on a performance-efficiency score weighted by a trade-off parameter $\alpha$. During inference, each query is embedded and mapped to its top-$p$ nearest clusters. The model with the highest performance-efficiency score aggregated over those clusters is selected to generate the response.
In our experiments, the Avengers-Pro consists of 8 models from 4 families: GPT-5-chat, GPT-5-medium, Claude-4.1-opus, Claude-4-sonnet, Gemini-2.5-pro, Gemini-2.5-flash, Qwen3-235B-A22B-thinking-2507, and Qwen3-235B-A22B-2507. We evaluate the Avengers-Pro on 6 challenging benchmarks: GPQA-Diamond rein2024gpqa , Human’s Last Exam phan2025humanity , HealthBench arora2025healthbench , ARC-AGI chollet2024arc , SimpleQA wei2024measuring , LiveCodeBench jain2024livecodebench , and $\tau$2-bench barres2025tau . We find that compared to the strongest single model GPT-5-medium (average accuracy: $62.25\%$, cost: \$47.96\\$47.96), the Avengers-Pro can attain $7\%$ performance gain with a comparable cost (average accuracy: $66.66$, cost: \$47.13\\$47.13), and cut $27\%$ cost with a comparable performance (average accuracy: $62.66$, cost: \$35.05\\$35.05). By varying the trade-off parameter $\alpha$, the Avengers-Pro achieves an even more favorable balance between performance and efficiency. For example, to reach 90% of GPT-5-medium’s performance—a level comparable to Gemini-2.5-pro—the Avengers-Pro reduces cost by 63% relative to GPT-5-medium and by 81% relative to Gemini-2.5-pro. Furthermore, we observe that the Avengers-Pro achieves a Pareto frontier: for any fixed cost, it consistently delivers the highest performance among all models at that expenditure. Conversely, for any fixed performance target, it provides the lowest cost compared to other models attaining the same accuracy.
2 Routing for Performance-Efficiency Trade-off
The Avengers-Pro ensembles a set of heterogeneous LLMs of varying capabilities and efficiencies with a router. Appropriate routing depends on an accurate understanding of each model’s capability and efficiency across different types of tasks or queries. To build this understanding, the router requires a set $\mathcal{D}$ of labeled query–answer pairs. Each query $d\in\mathcal{D}$ is first encoded into a semantic vector using a text embedding model. These embeddings are then grouped into $k$ clusters using a clustering algorithm, producing a set $\mathcal{C}=\{c_{1},\dots,c_{k}\}$, where each cluster represents a semantically coherent query type.
Let $\mathcal{M}$ denote the set of models in our system. We evaluate each model $i\in\mathcal{M}$ on $\mathcal{D}$, measures its performance and efficiency within each cluster. Let $\mathbf{p}^{i}=[p_{1}^{i},\dots,p_{k}^{j}]^{\top}$ be a cluster-wise performance profile for model $i$, where $p_{j}^{i}$ denotes model $i$’s accuracy on queries within cluster $c_{j}$. Similarly, let $\mathbf{q}^{i}=[q_{1}^{i},\dots,q_{k}^{j}]^{\top}$ be a cluster-wise efficiency profile for model $i$, where $q_{j}^{i}$ denotes model $i$’s efficiency on queries within cluster $c_{j}$. We measure the efficiency in terms of cost such that $q_{j}^{i}$ denotes the total cost incurred by model $i$ to answer all queries within cluster $c_{j}$.
We calculate the cluster-wise performance-efficiency score $x_{j}^{i}$ for model $i$ on $c_{j}$ by
$x_{j}^{i}=\alpha\,\tilde{p}_{j}^{i}+(1-\alpha)\,(1-\tilde{q}_{j}^{i}),$
where $\alpha\in[0,1]$ controls the trade-off between performance and efficiency, and $\tilde{p}_{j}^{i}$ and $\tilde{q}_{j}^{i}$ are the normalized values of $p_{j}^{i}$ and $q_{j}^{i}$. The normalization is given by
$\tilde{p}_{j}^{i}=\frac{p_{j}^{i}-p_{j}^{\min}}{p_{j}^{\max}-p_{j}^{\min}},\quad\tilde{q}_{j}^{i}=\frac{q_{j}^{i}-q_{j}^{\min}}{q_{j}^{\max}-q_{j}^{\min}},$
where $p_{j}^{\min}$ and $p_{j}^{\max}$ (or $q_{j}^{\min}$ and $q_{j}^{\max}$) denote the minimum and maximum performance (or cost) among all models for cluster $j$.
During inference, an incoming query is encoded with the text embedding model, and is assigned to the top-$p$ nearest cluster(s) in the embedding space. For each model $i\in\mathcal{M}$, we sum up its cluster-wise performance-efficiency scores over those top-$p$ clusters. The model with the highest sum of those scores is selected to generate the response.
3 Experiments
Our experiments compare the performance and efficiency of Avengers-Pro against leading single models.
3.1 Experimental Settings
Models
We consider 8 leading models, which vary in capability and efficiency, as follows:
We access these models through the OpenRouter API [^2] https://openrouter.ai/, as its standardized interface simplifies the process of running identical experiments across multiple models. The pricing for these models is detailed in Table 1. Prices for the Qwen3 family may vary across providers; throughout this paper we report the prices listed by OpenRouter.
Benchmarks
We consider 6 challenging benchmarks, as summarized in Table 2, covering advanced reasoning and general knowledge:
  • GPQA-Diamond rein2024gpqa : A graduate-level google-proof Q&A benchmark.
  • Human’s Last Exam (HLE) phan2025humanity : A frontier multi-modal benchmark of closed-ended academic questions. In this study, we use the text-only setting without custom patches, tool use, or retrieval during evaluation. For efficiency and reproducibility, we use the first 500 questions from the released pool and report accuracy under the official evaluation protocol.
  • ARC-AGI chollet2024arc : A benchmark focused on fluid intelligence, testing the ability to reason and solve novel problems. We use the first 200 questions from the released pool and report accuracy under the official evaluation protocol.
  • SimpleQA wei2024measuring : A factuality benchmark for short, fact-seeking questions. We use the official implementation with the default configuration and report accuracy under the official scoring. We evaluate on a subset of 500 examples uniformly sampled from the released dataset.
  • LiveCodeBench jain2024livecodebench : A dynamic, contamination-controlled coding benchmark that continuously ingests newly released problems. We evaluate on the latest public release (v6) using the official implementation and evaluation harness with the default configuration, without custom patches or post-processing.
  • $\tau^{2}$-bench barres2025tau : A controlled testbed for agents that must reason effectively and guide user actions.
For all benchmarks, we use the official repositories/implementations with their recommended parameter settings.
Implementation Details
We use k-means clustering with $k=60$ clusters. Each query is encoded by the Qwen3-embedding-8B zhang2025qwen3embeddingadvancingtext into a 4,096-dimensional semantic vector. Following common practice in routing routerdc ; zhuang2024embedllmlearningcompactrepresentations ; zhang2025avengers , we randomly split the data: 70% is used to fit the clustering model and estimate per-cluster statistics, and the remaining 30% is reserved for routing and evaluation. At inference time, we compute the embedding of the incoming query and retrieve the top-$p$ nearest clusters ($p=4$) in the embedding space. For each model $i$, we then sum its cluster-wise cost–capability scores $q_{j}^{i}$ over these three clusters and select the model with the highest total to generate the response.
Table 1: Model cost information (OpenRouter).
Model Input Price Output Price
(\$/1M tokens) (\$/1M tokens)
Gemini-2.5-flash 0.30 2.50
Gemini-2.5-Pro 1.25 10
Claude-4.1-opus 15 75
Claude-4-sonnet 3 15
GPT-5-chat 1.25 10
GPT-5-medium 1.25 10
Qwen3-235B-A22B-25074 $\approx$0.13 $\approx$0.6
Qwen3-235B-A22B-thinking-2507 $\approx$0.13 $\approx$0.6
Table 2: Benchmark information.
Dataset Metrics Size
ARC-AGI-v1 chollet2024arc pass@1 200
GPQA-Diamond rein2024gpqa pass@1 198
HLE phan2025humanity pass@1 500
LiveCodeBench-v6 jain2024livecodebench pass@1 1,055
$\tau^{2}$-bench barres2025tau pass@1 150
SimpleQA wei2024measuring pass@1 500
Total 2,603
3.2 Results and Analysis
We present the comparisons of Avengers-Pro and single models in terms of performance and efficiency in Table 3. We show how the trade-off parameter $\alpha$ affects the performance and efficiency in Figure 2. We show the proportion of model usage by Avengers-Pro in Figure 3.
Avengers-Pro outperforms top single models.
Of the eight single models evaluated, GPT-5-medium demonstrates the highest average accuracy ($62.25\%$) across the six benchmarks. This is followed by Gemini-2.5-pro ($56.08\%$) and Qwen3-thinking ($48.11\%$), respectively. The Avengers-Pro surpasses the performance of all individual models with a sufficiently large value of $\alpha$, prioritizing performance over efficiency. Specifically, its average accuracy is up to $66.66\%$ with $\alpha=1.0$, which is $7\%$ higher compared to GPT-5-medium and $19\%$ higher compared to Gemini-2.5-pro.
Avengers-Pro achieves a superior performance-efficiency trade-off.
At a performance level comparable to the strongest single model GPT-5-medium, Avengers-Pro ($\alpha=0.53$) incurs significantly lower costs, resulting in a cost reduction of $27\%$. Similarly, at a $90\%$ performance level of GPT-5-medium, the Avengers-Pro ($\alpha=0.39$) cuts cost by $63\%$. At a performance level comparable to the second-strongest single model Gemini-2.5-pro, it ($\alpha=0.39$) reduces cost by $81\%$. At a performance level comparable to Cluade-4.1-opus, it ($\alpha=0.25$) achieves a cost reduction of $92\%$. Moreover, as shown in Figure 1A, the Avengers-Pro achieves a Pareto frontier—no single model can simultaneously deliver higher performance and greater efficiency than Avengers-Pro. In other words, Avengers-Pro offers the highest performance for any given cost and the lowest cost for any given level of performance.
Table 3: The performance and efficiency of Avengers-Pro vs. single models. Note that GPT-5-chat has no score on the $\tau^{2}$-bench benchmark because this model does not support tool calling. Bold indicates the best performance of a given benchmark. With $\alpha=0.1$, Avengers-Pro, surpasses GPT-5-medium in average accuracy with a $7\%$ performance gain. With $\alpha=0.53$, it matches GPT-5-medium’s average accuracy, while cutting the cost by $27\%$. With $\alpha=0.39$, it reaches 90% of GPT-5-medium’s performance at a $63\%$ lower cost.
Setting ARC-AGI GPQA-Diamond HLE LiveCodeBench SimpleQA $\tau^{2}$-bench Avg. A Cost
Gemini-2.5-flash 9.62 21.72 7.20 62.84 28.99 36.67 27.84 \$7.10
Gemini-2.5-pro 33.08 84.85 23.09 78.67 54.80 62.00 56.08 \$94.87
Claude-4.1-opus 22.12 74.24 6.41 64.07 31.00 74.00 45.31 \$117.40
Claude-4-sonnet 16.15 68.69 4.60 59.05 15.00 64.00 37.92 \$25.35
Qwen3 9.22 58.59 9.22 66.26 53.00 53.33 41.60 \$2.73
Qwen3-thinking 19.23 80.81 12.68 77.99 44.60 53.33 48.11 \$13.99
GPT-5-chat 6.73 73.73 7.80 63.60 40.20 - 38.41 \$4.04
GPT-5-medium 44.42 84.85 26.20 88.44 47.60 82.00 62.25 \$47.96
Avengers Pro ($\alpha=0$) 15.33 58.67 10.13 66.94 46.27 0.00 32.89 \$1.08
Avengers Pro ($\alpha=0.25$)1 29.33 67.00 10.00 76.53 53.60 72.89 51.56 \$9.69
Avengers Pro ($\alpha=0.39$)2 29.33 78.67 12.67 84.79 55.07 76.89 56.24 \$17.81
Avengers Pro ($\alpha=0.53$)3 51.67 80.00 25.46 87.45 54.93 76.44 62.66 \$35.05
Avengers Pro ($\alpha=0.8$) 59.67 81.00 27.60 89.34 56.93 78.22 65.46 \$44.65
Avengers Pro ($\alpha=1$) 59.67 85.67 28.67 89.59 56.40 80.00 66.66 \$47.13
Effects of the trade-off parameter
As shown in Figure 2, we gradually increase the trade-off parameter $\alpha$, placing more weight on performance over efficiency. As $\alpha$ increases, the average accuracy increases rapidly for small $\alpha$ and then plateaus near $\alpha\!\approx\!0.6$. On the other hand, as $\alpha$ increases, cost remains low until about $\alpha\!\approx\!0.4$ before rising sharply. These trends reveal two elbows (around $0.4$ and $0.6$) that offer favorable trade-offs.
Figure 2: Effects of the trade-off parameter $\alpha$ on the performance and efficiency. A greater value of $\alpha$ prioritizes performance over efficiency. The increase in performance is usually accompanied the increase in cost.
Figure 3: Proportion of model usage, given various trade-off parameters $\alpha$. When $\alpha$ is low, Avengers-Pro tend to route queries to Qwen3 and Qwen3-thinking. With a greater value of $\alpha$, Avengers-Pro favors GPT5-medium and Qwen3-thinking.
Proportion of model usage
As shown in Figure 3, when $\alpha$ is low, Avengers-Pro tends to favor the Qwen3 and Qwen3-thinking model, routing a great proportion of queries to these two models with a low unit price. As $\alpha$ increases, the usage of GPT-5-medium rises rapidly; concurrently, the usage of Gemini-2.5-pro and Claude-opus-4.1, which excel at complex reasoning but have a higher unit price, also increases.
4 Conclusions
In this work, we introduce Avengers-Pro, a test-time routing framework integrating different LLMs to optimize the trade-off between performance and efficiency. By dynamically selecting exactly one model for each incoming query, Avengers-Pro optimizes both cost and accuracy. Our experiments involving 8 leading LLMs and 6 challenging benchmarks demonstrate that Avengers-Pro can surpass the strongest single model, GPT-5-medium, by up to 7% in accuracy and match its performance at a 27% lower expense. Moreover, Avengers-Pro achieves a Pareto frontier, consistently delivering the best performance on any given budget and the lowest cost given any performance target. Our results highlight the significant potential of an intelligent test-time routing framework in creating more powerful, efficient, and scalable LLM systems.
Acknowledgments
This work was supported by a locally commissioned task from the Shanghai Municipal Government. The authors thank Dr. Qiaosheng Zhang for helpful discussions, and collaborators who contributed meaningfully to our earlier project Avengers zhang2025avengers .